OpenStudy (anonymous):

prove that \[\cos(\cos(\cos......\cos(\theta ))))))...)=1 \]

OpenStudy (anonymous):

no...the question is right...

OpenStudy (anonymous):

Use the cosine expansion? \[\cos(x)=\sum_{k=0}^{\infty}\frac{(-1)^kx^{2k}}{(2k)!}\]

OpenStudy (anonymous):

in the problem i have used infinite times cos(cos.... so can u solve by that way...!!!!!

OpenStudy (anonymous):

how can it be infinite if theres a starting point and an ending point? Clearly there is one cos on the outside and one on the very very inside. If you treat it as an infinite amount im getting 1 = 0, which is why i thought it was finite instead.

OpenStudy (anonymous):

this problem will be solved analytically so go for that don't think of such big formulas.... if u have done the cosine problem then do the same using sine and prove that in that case answer will be zero.

OpenStudy (anonymous):

how do u find 1=0 can u proveit???

OpenStudy (anonymous):

*prove it

OpenStudy (anonymous):

have u done the previous problem???

OpenStudy (anonymous):

If it was infinite, i was doing this: \[\cos(\cos(\cos(\ldots \cos(\theta)\ldots)))) = 1 \] take arccos of both sides: \[\cos(\cos(\cos(\ldots \cos(\theta)\ldots))) = \cos^{-1}(1)\] If there were an infinite amount of cosines, then the left hand side is the same as the very beginning of the problem, and it equals 1: \[1 = \cos^{-1}(1) \iff 1 = 0\] Thats why i assumed it wasnt infinite, but finite.

OpenStudy (anonymous):

i havent done this problem before, but ive seen problems like: \[x^{x^{x^{x^{\ldots}}}} = 2\] which uses a similar technique.

OpenStudy (anonymous):

abywhos, i dont believe there is an infinite amount of cosines.

OpenStudy (anonymous):

anywhos*

OpenStudy (anonymous):

does anyone agree/disagree with me? any opinions on the matter? lolol

OpenStudy (zarkon):

the problem as stated is not true

OpenStudy (anonymous):

right, i think that as well.

OpenStudy (anonymous):

okay....wait...in case of sine how would you proceed if i ask to find the value of that function.

OpenStudy (anonymous):

why have you taken theta in the range .

OpenStudy (zarkon):

what do you mean by 'in case of sine how would you proceed if i ask to find the value of that function'

OpenStudy (anonymous):

find the value of sin(sin......sin(a))))...)

OpenStudy (zarkon):

0

OpenStudy (anonymous):

how ??

OpenStudy (anonymous):

any reason...?

OpenStudy (zarkon):

show that after the first application of sine the sequence is monotone. it will be obviously bounded. Monotone bounded sequences converge.

OpenStudy (anonymous):

exactly and by the same rule when the arguments become zero then cosine gives 1.

OpenStudy (anonymous):

do u agree with me?

OpenStudy (zarkon):

no

OpenStudy (zarkon):

the infinite regression of cosines give an answer other than 1

OpenStudy (anonymous):

why not ??

OpenStudy (zarkon):

why don't you try a few values of theta to see for yourself

OpenStudy (zarkon):

it will converge to a number that is approx ... 0.73908513321515

OpenStudy (zarkon):

the solution to the equation \[\cos(x)=x\]

OpenStudy (anonymous):

yes that's true ...i made a mistake...actually i compared with the sine one. you are right also the Gauss iteration formula also gives a solution like that of u

OpenStudy (anonymous):

okay its very nice to deal with you. hope we will meet again thanks both of you guys ...good job.